find the mean and standard deviation of a random variable with pdf f(x)=e^-x ,x>=0
1
Expert's answer
2020-05-11T18:28:18-0400

mean=E(x)="\int _0^{\infty \:} \:xe^{-x}dx=-e^{-x}x-e^{-x}" |"\infty" 0=0-(-1)=1

E(x2)="\int _0^{\infty \:}x^2e^{-x}dx" ="-e^{-x}x^2+2\left(-e^{-x}x-e^{-x}\right)" |∞ 0=0-(-2)=2

Var=E(x2)-E(x)2=4-1=3

SD="\sqrt {Var}" = 1.73